OpenStudy (anonymous):

how to graph : y=2cos½(x+π/2)

OpenStudy (luckycoins888):

ill help

OpenStudy (luckycoins888):

y=2sin(x-pi/2)find the amplitude, phas shift, and period for the graph of each function. y=cos(2x-pi/4) y=-4sin(2x/3+pi/6) y= 5/4cos(3x-2pi) y= 2 tan(2x-pi/4) .. standard forms used: y=Asin(Bx-C),A=amplitude, Period=2pi/B, Phase shift=C/B y=Acos(Bx-C),A=amplitude, Period=2pi/B, Phase shift=C/B y=tan(Bx-C),A=amplitude, Period=pi/B, Phase shift=C/B .. y=2sin(x-pi/2) B=1 Amplitude=2 Period=2pi/1=2pi x-pi/2=0 x=p/2 Phase Shift=(pi/2)to the right .. y=cos(2x-pi/4)= B=2 Amplitude=1 Period=2pi/2=pi 2x-pi/4=0 2x=pi/4 x=pi/8

OpenStudy (luckycoins888):

medal plz